OpenStudy (anonymous):

Mario decided to invest his $620 tax refund rather than spending it. He found a bank that would pay him 4% interest, compounded quarterly. Mario deposits the entire $620 refund and does not deposit or withdraw any other amount. a. Write an equation that models the growth of the investment. b. How many years will it take for the initial investment to double?

OpenStudy (anonymous):

\[A=P(1+\frac{r}{n})^{nt}\] \[A=620(1+\frac{.04}{4})^{4t}\]

OpenStudy (anonymous):

\[A=P(1+\frac{r}{n})^{nt}\] \[(1+\frac{r}{n})^{nt}=\frac{A}{P}\] \[ntlog(1+\frac{r}{n})=\log(\frac{A}{P})\] \[t=\frac{\log(\frac{A}{P})}{nlog(1+\frac{r}{n})}\] In your case: \[t=\frac{\log2}{4\log(1+\frac{.04}{4})}\] t=17.42 years

OpenStudy (anonymous):

Hi what does the nlog and log stands for???